ns ie tound in c?C?1C? Q2. How many neutrons are found in 12C? 13C?C? 12 C13 Q3. How many electrons are found in-C? 2. ? 13C? Q4. a) W ions (charged particles, shown by-or +sign in its symbol)? hich kind of subatomic particle (proton, neutron, or electron) is responsible for making atoms into Protons and neutons b) In terms of subatomic particles, explain how 13C got a negative charge compared to c (neutral, no charge). c) In terms of subatomic particles, explain why Na has a positive charge. Q5. Based on the model a) What do all carbon atoms and ions have in common? tome and ions have in common?
media%2F29d%2F29d05ef5-20ff-4ea6-a219-83
media%2F600%2F6007a514-b61e-417c-b584-34
Answer #1

2)
Atomic number of C = 6

mass number of 12C = 12
USE:
number of neutron = mass number - atomic number
= 12 - 6
= 6

mass number of 13C = 13
USE:
number of neutron = mass number - atomic number
= 13 - 6
= 7


mass number of 14C = 14
USE:
number of neutron = mass number - atomic number
= 14 - 6
= 8

So, answer will be 6, 7 and 8

I am allowed to answer only 1 question at a time